# DEXTER: Diffusion-Guided EXplanations with TExtual Reasoning for Vision Models

### v1

@@ 1 Introduction
This question is crucial as AI systems are increasingly deployed in high-stakes applications, where interpretability and trust are as important as accuracy. However, the lack of transparency in model reasoning, often exacerbated by the reliance on spurious correlations—irrelevant features that disproportionately influence predictions— undermines confidence in these systems, making decisions difficult to justify. For instance, ImageNet-trained classifiers have been shown to favor background textures or lighting conditions over intrinsic object properties [41, 9]. Addressing these issues requires explainability techniques that extend beyond local attribution, offering a global perspective on a model’s reasoning patterns, biases, and learned representations.
Existing methods for model explainability, such as GradCAM [39] and Integrated Gradients [43] provide local explanations and require data. These methods focus on analyzing individual predictions by attributing importance to specific pixels or regions in an image. While these methods are extremely useful for highlighting areas of interest, they do not offer a global understanding of model behavior. In contrast, activation maximization techniques [26, 27, 28] have been instrumental in globally visualizing the features learned by neural network, but the generated images are often abstract and challenging to interpret. While existing methods offer useful insights into model behavior and feature representations, they often fall short of capturing the global reasoning patterns and biases behind predictions. In particular, visual explanations can be hard to interpret and may lack the ability to convey high-level reasoning or reveal subtle spurious correlations—issues that are often better addressed through complementary textual explanations [18, 6, 16, 3].
Text-based explanations offer an accessible complement to visual methods [11], but they often lack global perspective or clarity. Recent work [16, 2, 10], including Natural Language Explanation (NLE) methods [30, 37, 15], combines visual and textual cues to reveal model biases, but typically relies on labeled data, annotations, or pretrained vision-language mappings [2].
@@
We evaluate DEXTER across three tasks: (1) activation maximization to reveal model-prioritized features, (2) slice discovery to detect underperforming subpopulations, and (3) bias explanation to identify spurious correlations. Experiments use SalientImageNet [41], Waterbirds [36], CelebA [21], and FairFaces [14]—datasets widely used in fairness and interpretability research. Across tasks, DEXTER demonstrates strong performance, offering meaningful insights into model behavior.
@@ 2 Related work
Activation Maximization (AM) interprets neural networks by generating inputs that maximize neuron activations [7]. Early methods often produced unrealistic, uninterpretable images [40, 25], later improved through regularization [23] and generative priors [26]. DiffExplainer [31] introduced diffusion-based image generation guided by soft prompts, which improve realism but lack semantic transparency. This work points toward token-wise optimization, but this direction remains underexplored. In contrast, DEXTER builds on this line by replacing soft prompts with discrete (hard) tokens, which are interpretable and enable both visual and textual global explanations. Although once central to model interpretability, AM has seen limited progress in recent years, partly because the resulting images—while optimized for neuron activation—are often difficult to interpret semantically, making it challenging to draw clear, causal insights about model behavior. The field has largely shifted toward attribution methods—e.g., GradCAM [39], Integrated Gradients [43], which are computationally efficient and provide intuitive saliency maps, though they are inherently local and data-dependent. Other efforts in feature visualization [52] offer insights into internal representations but require training data and remain focused on visual output. DEXTER revives AM by leveraging modern diffusion models and discrete prompt optimization to produce class-level, multimodal explanations in a fully data-free setting.
@@
Explanation of visual classifiers. Textual and Multimodal Explanations seek to complement visual saliency with natural language justifications. Approaches like Multimodal Explanations [29], LIMEx [44], and post-hoc counterfactuals [48] align vision and language to explain model decisions. However, these methods often rely on ground-truth labels or annotated datasets and are limited to instance-level explanations. In contrast, DEXTER provides global, class-wise explanations without supervision or access to data.
@@
Natural Language Explanation (NLE) methods aim to generate human-readable justifications, typically using VQA-style benchmarks [30, 37, 15]. Recent approaches integrate vision and language into unified architectures [38, 37], but remain supervised and local. DEXTER differs by producing global textual reports through unsupervised classifier probing—without labels, data, or task-specific fine-tuning.
@@
Slice Discovery identifies dataset subpopulations where a model underperforms, offering a targeted way to reveal and explain systematic failures in classifier behavior. Traditional methods analyze embeddings, gradients, or misclassifications [1, 24, 42, 51], while recent approaches leverage CLIP’s joint text-visual embeddings [8, 12, 50] or use LLMs to generate captions and extract keywords [46, 16]. Bias-to-Text (B2T)[16] extracts pseudo-bias labels from misclassified image captions, and LADDER[10] generates bias hypotheses from low-confidence predictions, clustering samples via LLM-derived pseudo-attributes. Unlike these data-dependent methods, DEXTER discovers and explains biases in a fully data-free manner, relying only on the classifier’s internal behavior.
@@ 3 Method
DEXTER’s framework, shown in Fig. 1, integrates three key components: a text pipeline for optimizing prompts, a vision pipeline for the image generation process, and a reasoning module using a vision-language model (VLM). DEXTER begins by optimizing a soft prompt to condition a BERT model [5] to fill in masked tokens in a predefined sentence. The resulting prompt guides the stable diffusion process to generate images that maximize the activation of a set of target neurons (e.g. classification heads) in a given visual classifier. The generated images are then analyzed by the VLM, which reasons across multiple images to provide coherent, human-readable textual explanations of the model’s decision-making process.
@@ 3.1 Text pipeline
The text pipeline has the goal of optimizing a textual prompt to suitably condition the diffusion model. We pose prompt generation as a masked language modeling task, and employ a pretrained and frozen BERT model for this purpose. The structure of the textual prompt to be produced is fully customizable, and can be controlled by combining portions of fixed text with a set of mask tokens. For the sake of clarity and without loss of generality, we will assume that the textual prompt has the structure of a sequence $\mathbf{t}=\left[\mathbf{t}_{\text{fixed}},m_{1},m_{2},\dots,m_{N}\right]$, where $\mathbf{t}_{\text{fixed}}$ is the portion of fixed text and all $m_{i}$ are set to BERT’s [MASK]token. Let $\mathbf{t}_{\text{emb}}$ be the embedding of $\mathbf{t}$, including positional encoding. In order to alter BERT’s behavior, which would naturally tend to replace $m_{i}$ with the most likely tokens based on its pretraining, we also prepend to the input sequence a learnable soft prompt $\mathbf{p}\in\mathbb{R}^{P\times d}$, consisting of a sequence of $P$ vectors, with $d$ being the dimensionality of BERT’s embedding space [17]. The full input sequence to BERT is thus $\left[\mathbf{p},\mathbf{t}_{\text{emb}}\right]$. We read out the logits corresponding to the masked tokens, i.e., $\left[\mathbf{l}_{1},\dots,\mathbf{l}_{N}\right]$, where each $\mathbf{l}_{i}\in\mathbb{R}^{V}$ and $V$ is BERT’s vocabulary size. Each logit vector $\mathbf{l}_{i}$ is mapped to a differentiable one-hot vector $\mathbf{o}_{i}\in\left\{0,1\right\}^{V}$, $\sum o_{i,j}=1$, through a Gumbel-softmax [22, 13] (with temperature $\tau=1$), from which the corresponding predicted token $\hat{t}_{i}\in\left\{1,\dots,V\right\}$ is retrieved. The resulting text prompt can be recovered as $\left[\mathbf{t}_{\text{fixed}},\hat{t}_{1},\dots,\hat{t}_{N}\right]$.
@@
At this point, a practical problem arises, in that standard implementations of diffusion models (e.g., Stable Diffusion [35]) CLIP’s text encoder [32] is employed to embed textual prompts into a conditioning vector. Unfortunately, BERT’s and CLIP’s vocabulary overlap only partially. To address this issue, we employ a translation matrix $\mathbf{M}\in\{0,1\}^{V\times W}$ to map each one-hot vector $\mathbf{o}_{i}$ to its corresponding representation in CLIP’s vocabulary, of size $W$. In $\mathbf{M}$, each row contains a single 1, indicating the index of the corresponding token in the CLIP vocabulary. Thus, given an original one-hot vector $\mathbf{o}_{i}$ provided by BERT, we can translate it into its CLIP equivalent through $\mathbf{o}_{i}^{(C)}=\mathbf{o}_{i}\mathbf{M}$, indexing a token $\hat{t}_{i}^{\text{(C)}}$ in CLIP’s vocabulary. For unassigned BERT tokens, the corresponding rows in $\mathbf{M}$ are entirely zero. As a result of this design, the model learns to avoid predicting BERT tokens that do not have a valid mapping in CLIP, as they correspond to a zero indexing vector $\mathbf{o}_{i}^{(C)}$, which would provide a meaningless representation when multiplied by the CLIP embedding look-up table.
@@ 4.1 Datasets
SalientImageNet [41] is used to evaluate both visual explanations and activation maximization. It is a curated subset of ImageNet designed to analyze model reasoning through object and context annotations. For each class, the top-5 neural features—highly activated units in the penultimate layer—are annotated as either spurious or core, based on whether they reflect incidental or meaningful correlations with the target label. This enables a fine-grained assessment of DEXTER’s ability to highlight robust, semantically grounded features in both explanation and feature synthesis settings.
Waterbirds[36] and CelebA[21] serve as benchmarks for bias discovery. Waterbirds introduces spurious correlations between bird species and backgrounds (e.g., land vs. water), providing a controlled environment for evaluating slice discovery and DEXTER’s ability to surface underperforming subpopulations. CelebA offers over 200,000 annotated face images with 40 binary attributes (e.g., gender, hairstyle, glasses), allowing us to assess how DEXTER identifies biases in classifier behavior related to demographic and semantic features.
FairFaces [14] is finally used for bias text explanation, focusing on the articulation of systematic spurious correlations in facial classification tasks. With over 100,000 images balanced across seven demographic groups, it enables a rigorous evaluation of how DEXTER captures and communicates bias in decision-making across diverse populations.
@@ 4.2.1 Visual Explanations
Fig. 3 presents an example of visual and textual explanations generated by DEXTER for RobustResNet50 [41], focusing on the 5 most active features in the penultimate layer for the dog sled class, which are all categorized as spurious in SalientImageNet. DEXTER-generated images align more effectively with the classifier’s attended regions compared to those generated by DiffExplainer. Furthermore, DEXTER provides textual descriptions that clearly explain the semantics of each neural feature, offering an interpretable account of the classifier’s reasoning. This approach enhances interpretability and enables users to either identify spurious correlations, such as the ones in the dog sled class where none of the five most important features actually include a sled, or confirm the classifier’s reliance on meaningful, task-relevant attributes.
For quantitative evaluation, we compare DEXTER to DiffExplainer using CLIP-IQA and Semantic CLIP-IQA [45] (Appendix D). DEXTER achieves higher scores (0.94 ± 0.03 / 0.96 ± 0.03) than DiffExplainer (0.89 ± 0.09 / 0.89 ± 0.09), indicating better semantic alignment and consistency. Additional comparisons with GAN-based methods [47, 23, 26] further confirm the advantage of diffusion-based explanations.
@@ 4.2.2 Slice discovery and debiasing
The goal of slice discovery is to identify subgroups of data where the model exhibits worse performance compared to the rest of the dataset. We assess slice discovery and debiasing performance using CelebA and Waterbirds datasets. In the former, the task is “blonde”/“non-blonde” classification, with the “blonde” class including a low proportion of men (representing a slice). The Waterbirds dataset tackles “waterbird“/“landbird” classification, and is built so that a portion of waterbird images feature a land background, and vice versa, thus introducing a slice per class. We perform slice discovery with DEXTER by first identifying words that maximize the activation of a given class. Specifically, we leverage DEXTER’s optimization process to obtain several descriptive words for each class. Following this step, as done by Kim et al. [16], we compute the CLIP similarity between the discovered words and the images in the dataset: images with high similarity are identified as belonging to a slice. We thus train a debiased classifier using Distributionally Robust Optimization (DRO) [36, 33] and evaluate the classification accuracy on the identified slices. We compare this approach with ERM [36], LfF [24], GEORGE [42], JTT [19], CNC [49], DRO [36], LADDER [10] and DRO-B2T [16]. As shown in Tab. 1, DEXTER outperforms state-of-the-art methods on worst-slice prediction for CelebA and achieves comparable results on Waterbirds, deriving descriptive words directly from the model without relying on training data. Additional details on the slice discovery task and results are given in Appendix E.

### v2

@@ 1 Introduction
This question is crucial as AI systems are increasingly deployed in high-stakes applications, where interpretability and trust are as important as accuracy. However, the lack of transparency in model reasoning, often exacerbated by the reliance on spurious correlations—irrelevant features that disproportionately influence predictions— undermines confidence in these systems, making decisions difficult to justify. For instance, ImageNet-trained classifiers have been shown to favor background textures or lighting conditions over intrinsic object properties [singla2022salient, geirhos2018imagenettrained]. Addressing these issues requires explainability techniques that extend beyond local attribution, offering a global perspective on a model’s reasoning patterns, biases, and learned representations.
Existing methods for model explainability, such as GradCAM [8237336] and Integrated Gradients [sundararajan2017axiomatic] provide local explanations and require data. These methods focus on analyzing individual predictions by attributing importance to specific pixels or regions in an image. While these methods are extremely useful for highlighting areas of interest, they do not offer a global understanding of model behavior. In contrast, activation maximization techniques [nguyen2016synthesizing, nguyen2017plug, olah2017feature] have been instrumental in globally visualizing the features learned by neural network, but the generated images are often abstract and challenging to interpret. While existing methods offer useful insights into model behavior and feature representations, they often fall short of capturing the global reasoning patterns and biases behind predictions. In particular, visual explanations can be hard to interpret and may lack the ability to convey high-level reasoning or reveal subtle spurious correlations—issues that are often better addressed through complementary textual explanations [lipton2018mythos, doshi2017towards, kim2024discovering, chen2022rex].
Text-based explanations offer an accessible complement to visual methods [hendricks2016generating], but they often lack global perspective or clarity. Recent work [kim2024discovering, asgari2024texplain, ghosh2025ladderlanguagedrivenslice], including Natural Language Explanation (NLE) methods [park2018multimodal, NLE_UniNLX, NLE_eViL], combines visual and textual cues to reveal model biases, but typically relies on labeled data, annotations, or pretrained vision-language mappings [asgari2024texplain].
@@
We evaluate DEXTER across three tasks: (1) activation maximization to reveal model-prioritized features, (2) slice discovery to detect underperforming subpopulations, and (3) bias explanation to identify spurious correlations. Experiments use SalientImageNet [singla2022salient], Waterbirds [sagawa2019distributionally], CelebA [liu2015faceattributes], and FairFaces [karkkainen2021fairface]—datasets widely used in fairness and interpretability research. Across tasks, DEXTER demonstrates strong performance, offering meaningful insights into model behavior.
@@ 2 Related work
Activation Maximization (AM) interprets neural networks by generating inputs that maximize neuron activations [erhan2009visualizing]. Early methods often produced unrealistic, uninterpretable images [simonyan2014deep, nguyen2015deep], later improved through regularization [mahendran2016visualizing] and generative priors [nguyen2016synthesizing]. DiffExplainer [pennisi2024diffexplainer] introduced diffusion-based image generation guided by soft prompts, which improve realism but lack semantic transparency. This work points toward token-wise optimization, but this direction remains underexplored. In contrast, DEXTER builds on this line by replacing soft prompts with discrete (hard) tokens, which are interpretable and enable both visual and textual global explanations. Although once central to model interpretability, AM has seen limited progress in recent years, partly because the resulting images—while optimized for neuron activation—are often difficult to interpret semantically, making it challenging to draw clear, causal insights about model behavior. The field has largely shifted toward attribution methods—e.g., GradCAM [8237336], Integrated Gradients [sundararajan2017axiomatic], which are computationally efficient and provide intuitive saliency maps, though they are inherently local and data-dependent. Other efforts in feature visualization [zimmermann2021featurevisualizations] offer insights into internal representations but require training data and remain focused on visual output. DEXTER revives AM by leveraging modern diffusion models and discrete prompt optimization to produce class-level, multimodal explanations in a fully data-free setting.
@@
Explanation of visual classifiers. Textual and Multimodal Explanations seek to complement visual saliency with natural language justifications. Approaches like Multimodal Explanations [park2018multimodal, NLE_UniNLX], and post-hoc counterfactuals [asgari2024texplain] align vision and language to explain model decisions. However, these methods often rely on ground-truth labels or annotated datasets and or limited to instance-level explanations. In contrast, DEXTER provides global, class-wise explanations without supervision or access to data.
@@
Natural Language Explanation (NLE) methods aim to generate human-readable justifications, typically using VQA-style benchmarks [park2018multimodal, NLE_UniNLX, NLE_eViL]. Recent approaches integrate vision and language into unified architectures [NLE_NLXGPT, NLE_UniNLX], but remain supervised and local. DEXTER differs by producing global textual reports through unsupervised classifier probing—without labels, data, or task-specific fine-tuning.
@@
Slice Discovery identifies dataset subpopulations where a model underperforms, offering a targeted way to reveal and explain systematic failures in classifier behavior. Traditional methods analyze embeddings, gradients, or misclassifications [agarwal2022estimating, nam2020learning, sohoni2020no, zhao2023combating], while recent approaches leverage CLIP’s joint text-visual embeddings [eyubogludomino, jaindistilling, zhang2022drml] or use LLMs to generate captions and extract keywords [wilesdiscovering, kim2024discovering]. Bias-to-Text (B2T)[kim2024discovering] extracts pseudo-bias labels from misclassified image captions, and LADDER[ghosh2025ladderlanguagedrivenslice] generates bias hypotheses from low-confidence predictions, clustering samples via LLM-derived pseudo-attributes. Unlike these data-dependent methods, DEXTER discovers and explains biases in a fully data-free manner, relying only on the classifier’s internal behavior.
@@ 3 Method
DEXTER’s framework, shown in Fig. 1, integrates three key components: a text pipeline for optimizing prompts, a vision pipeline for the image generation process, and a reasoning module using a vision-language model (VLM). DEXTER begins by optimizing a soft prompt to condition a BERT model [devlin2018bert] to fill in masked tokens in a predefined sentence. The resulting prompt guides the stable diffusion process to generate images that maximize the activation of a set of target neurons (e.g. classification heads) in a given visual classifier. The generated images are then analyzed by the VLM, which reasons across multiple images to provide coherent, human-readable textual explanations of the model’s decision-making process.
@@ 3.1 Text pipeline
The text pipeline has the goal of optimizing a textual prompt to suitably condition the diffusion model. We pose prompt generation as a masked language modeling task, and employ a pretrained and frozen BERT model for this purpose. The structure of the textual prompt to be produced is fully customizable, and can be controlled by combining portions of fixed text with a set of mask tokens. For the sake of clarity and without loss of generality, we will assume that the textual prompt has the structure of a sequence $\mathbf{t}=\left[\mathbf{t}_{\text{fixed}},m_{1},m_{2},\dots,m_{N}\right]$, where $\mathbf{t}_{\text{fixed}}$ is the portion of fixed text and all $m_{i}$ are set to BERT’s [MASK]token. Let $\mathbf{t}_{\text{emb}}$ be the embedding of $\mathbf{t}$, including positional encoding. In order to alter BERT’s behavior, which would naturally tend to replace $m_{i}$ with the most likely tokens based on its pretraining, we also prepend to the input sequence a learnable soft prompt $\mathbf{p}\in\mathbb{R}^{P\times d}$, consisting of a sequence of $P$ vectors, with $d$ being the dimensionality of BERT’s embedding space [lester2021power]. The full input sequence to BERT is thus $\left[\mathbf{p},\mathbf{t}_{\text{emb}}\right]$. We read out the logits corresponding to the masked tokens, i.e., $\left[\mathbf{l}_{1},\dots,\mathbf{l}_{N}\right]$, where each $\mathbf{l}_{i}\in\mathbb{R}^{V}$ and $V$ is BERT’s vocabulary size. Each logit vector $\mathbf{l}_{i}$ is mapped to a differentiable one-hot vector $\mathbf{o}_{i}\in\left\{0,1\right\}^{V}$, $\sum o_{i,j}=1$, through a Gumbel-softmax [maddison2017concrete, jang2017categorical] (with temperature $\tau=1$), from which the corresponding predicted token $\hat{t}_{i}\in\left\{1,\dots,V\right\}$ is retrieved. The resulting text prompt can be recovered as $\left[\mathbf{t}_{\text{fixed}},\hat{t}_{1},\dots,\hat{t}_{N}\right]$.
@@
At this point, a practical problem arises, in that standard implementations of diffusion models (e.g., Stable Diffusion [rombach2022high]) CLIP’s text encoder [radford2021learning] is employed to embed textual prompts into a conditioning vector. Unfortunately, BERT’s and CLIP’s vocabulary overlap only partially. To address this issue, we employ a translation matrix $\mathbf{M}\in\{0,1\}^{V\times W}$ to map each one-hot vector $\mathbf{o}_{i}$ to its corresponding representation in CLIP’s vocabulary, of size $W$. In $\mathbf{M}$, each row contains a single 1, indicating the index of the corresponding token in the CLIP vocabulary. Thus, given an original one-hot vector $\mathbf{o}_{i}$ provided by BERT, we can translate it into its CLIP equivalent through $\mathbf{o}_{i}^{(C)}=\mathbf{o}_{i}\mathbf{M}$, indexing a token $\hat{t}_{i}^{\text{(C)}}$ in CLIP’s vocabulary. For unassigned BERT tokens, the corresponding rows in $\mathbf{M}$ are entirely zero. As a result of this design, the model learns to avoid predicting BERT tokens that do not have a valid mapping in CLIP, as they correspond to a zero indexing vector $\mathbf{o}_{i}^{(C)}$, which would provide a meaningless representation when multiplied by the CLIP embedding look-up table.
@@ 4.1 Datasets
SalientImageNet [singla2022salient] is used to evaluate both visual explanations and activation maximization. It is a curated subset of ImageNet designed to analyze model reasoning through object and context annotations. For each class, the top-5 neural features—highly activated units in the penultimate layer—are annotated as either spurious or core, based on whether they reflect incidental or meaningful correlations with the target label. This enables a fine-grained assessment of DEXTER’s ability to highlight robust, semantically grounded features in both explanation and feature synthesis settings.
Waterbirds[sagawa2019distributionally] and CelebA[liu2015faceattributes] serve as benchmarks for bias discovery. Waterbirds introduces spurious correlations between bird species and backgrounds (e.g., land vs. water), providing a controlled environment for evaluating slice discovery and DEXTER’s ability to surface underperforming subpopulations. CelebA offers over 200,000 annotated face images with 40 binary attributes (e.g., gender, hairstyle, glasses), allowing us to assess how DEXTER identifies biases in classifier behavior related to demographic and semantic features.
FairFaces [karkkainen2021fairface] is finally used for bias text explanation, focusing on the articulation of systematic spurious correlations in facial classification tasks. With over 100,000 images balanced across seven demographic groups, it enables a rigorous evaluation of how DEXTER captures and communicates bias in decision-making across diverse populations.
@@ 4.2.1 Visual Explanations
Fig. 3 presents an example of visual and textual explanations generated by DEXTER for RobustResNet50 [singla2022salient], focusing on the 5 most active features in the penultimate layer for the dog sled class, which are all categorized as spurious in SalientImageNet. DEXTER-generated images align more effectively with the classifier’s attended regions compared to those generated by DiffExplainer. Furthermore, DEXTER provides textual descriptions that clearly explain the semantics of each neural feature, offering an interpretable account of the classifier’s reasoning. This approach enhances interpretability and enables users to either identify spurious correlations, such as the ones in the dog sled class where none of the five most important features actually include a sled, or confirm the classifier’s reliance on meaningful, task-relevant attributes.
For quantitative evaluation, we compare DEXTER to DiffExplainer using CLIP-IQA and Semantic CLIP-IQA [wang2022exploring] (Appendix D). DEXTER achieves higher scores (0.94 ± 0.03 / 0.96 ± 0.03) than DiffExplainer (0.89 ± 0.09 / 0.89 ± 0.09), indicating better semantic alignment and consistency. Additional comparisons with GAN-based methods [yosinski2015understanding, mahendran2016visualizing, nguyen2016synthesizing] further confirm the advantage of diffusion-based explanations.
@@ 4.2.2 Slice discovery and debiasing
The goal of slice discovery is to identify subgroups of data where the model exhibits worse performance compared to the rest of the dataset. We assess slice discovery and debiasing performance using CelebA and Waterbirds datasets. In the former, the task is “blonde”/“non-blonde” classification, with the “blonde” class including a low proportion of men (representing a slice). The Waterbirds dataset tackles “waterbird“/“landbird” classification, and is built so that a portion of waterbird images feature a land background, and vice versa, thus introducing a slice per class. We perform slice discovery with DEXTER by first identifying words that maximize the activation of a given class. Specifically, we leverage DEXTER’s optimization process to obtain several descriptive words for each class. Following this step, as done by kim2024discovering, we compute the CLIP similarity between the discovered words and the images in the dataset: images with high similarity are identified as belonging to a slice. We thus train a debiased classifier using Distributionally Robust Optimization (DRO) [sagawa2019distributionally, rahimian2019distributionally] and evaluate the classification accuracy on the identified slices. We compare this approach with ERM [sagawa2019distributionally], LfF [nam2020learning], GEORGE [sohoni2020no], JTT [liu2021just], CNC [zhang2022correct], DRO [sagawa2019distributionally], LADDER [ghosh2025ladderlanguagedrivenslice] and DRO-B2T [kim2024discovering]. As shown in Tab. 1, DEXTER outperforms state-of-the-art methods on worst-slice prediction for CelebA and achieves comparable results on Waterbirds, deriving descriptive words directly from the model without relying on training data. Additional details on the slice discovery task and results are given in Appendix E.